OpenStudy (anonymous):

Given the set of scores 5.4, 5.6, 5.7, 5.9, 6.0; find the variance. A ) 0.381 B ) None of these C ) 0.0456 D ) 0.145 E ) 0.214

Parth (parthkohli):

B) None of these, the variance is 0.057

OpenStudy (anonymous):

What is variance ? Could u explain how u solve @ParthKohli pls.

OpenStudy (anonymous):

and where s my solution?

Parth (parthkohli):

Actually I know mean, median,mode.. I don't know variance haaha

OpenStudy (amistre64):

\[var=\frac{\sum (x-\mu)^2}{N}\]

OpenStudy (amistre64):

u = 5.72 since we square the results; the order you subtract aint important 5.4-u = .32: 5.6-u = .12: 5.7-u = .02: 5.9-u = .18: 6.0-u = .28: squre the results and add em up 5.4-u = .32: .1024 5.6-u = .12: .0144 5.7-u = .02: .0004 5.9-u = .18: .0324 6.0-u = .28: .0784 ------- .2280/5 = .0456